# femtoPro: Real-time linear and nonlinear optics simulations

### v1

@@ 5 Graphical laser beam representation
Simulating linear and nonlinear optical phenomena leads to laser beams propagating through space with parameters as determined in Section 2, Section 3, and Section 4. If one only seeks numerical results, then the calculation is already complete because the obtained parameters provide a full characterization. Often it is desired, however, to represent the laser beams graphically without compromising the real-time simulation. The details of graphical visualization of laser beams will depend on the particular software implementation, which is beyond the scope of the present work. Instead, we discuss some general points that are relevant for any such three-dimensional modeling that is sufficiently fast and realistic.
5.1 Gaussian beam scaling
@@
In this section, we investigate general spatial properties of Gaussian beams. The question will be how the shapes of Gaussian laser beams are related to each other given arbitrary beam parameters. We will find that any Gaussian beam can be derived from the general shape of a “reference beam” by suitable scaling. Thus, it is not necessary to re-calculate, point by point, the spatial envelopes of a beam for any new set of beam parameters, but one can simply generate a Gaussian model once, as will be shown in Section 5.2, and then scale it accordingly.
Let us assume that we start with a suitable representation of a Gaussian “reference beam” that is given by the beam radius,
$\tilde{w}(z)=\tilde{w}_{0}\sqrt{1+\left(\frac{z}{\tilde{z}_{\text{R}}}\right)^{2}},$ (61)
as a function of the distance $z$ from the beam waist with a “Rayleigh length”
$\tilde{z}_{\text{R}}=\frac{\pi\tilde{w}_{0}^{2}}{\lambda M^{2}},$ (62)
for propagation along the $\hat{\mathbf{z}}$ direction, wavelength $\lambda$, and beam-quality parameter $M^{2}\geq 1$ (for Gaussian beams, $M^{2}=1$), with the waist $\tilde{w}_{0}$ located at $z=0$. The tilde symbolizes the “reference” character. We seek a new beam
$w(z)=w_{0}\sqrt{1+\left(\frac{z-z_{0}}{z_{\text{R}}}\right)^{2}}$ (63)
with a new Rayleigh length
$z_{\text{R}}=\frac{\pi w_{0}^{2}}{\lambda M^{2}}$ (64)
whose (potentially different) waist $w_{0}$ is located at position $z=z_{0}$, and we would like to obtain this beam as a suitably transformed version of the reference beam.
It is often convenient to treat Gaussian beam propagation using a complex-valued beam curvature radius $q$, as introduced in Eq. (LABEL:SIEqComplexRadiusOfCurvature) in the SI, that is defined as [45]
$\frac{1}{q(z)}=\frac{1}{R(z)}+i\frac{\lambda M^{2}}{\pi w^{2}(z)}$ (65)
at a position $z$ with relation to the wavefront curvature radius $R$. Beam propagation in free space follows [46, 47]
$q(z)=q_{0}+z$ (66)
for propagation by $z$, and $q_{0}=q(0)$ [see also Eq. (LABEL:SIEqPropagationFreeSpace) in the SI]. Optical elements such as lenses or mirrors will lead to a modification of $q$ [14].
According to Eq. (66), we have
$\tilde{q}(z)=\tilde{q}_{0}+z$ (67)
and
$\tilde{q}_{0}=-i\tilde{z}_{\text{R}}$ (68)
for the propagation law of the reference beam, and its complex curvature at its waist position, respectively. Assume we are given $q=\operatorname{Re}q+i\operatorname{Im}q$, with $\operatorname{Re}q$ and $\operatorname{Im}q$ providing the real and imaginary parts of the complex radius of curvature $q$ of the “new” beam, immediately after transfer through an optical element. Then the new beam waist is located at a position $z_{0}=-\operatorname{Re}q$, translated with respect to the current position at the optical element, and the transformed Rayleigh length is given by $z_{R}=-\operatorname{Im}q$.
Let us define the scaling factor
$b=\frac{z_{\text{R}}}{\tilde{z}_{\text{R}}}=-\frac{\operatorname{Im}q}{\tilde{z}_{\text{R}}}$ (69)
such that
$z_{\text{R}}=b\tilde{z}_{\text{R}}$ (70)
and
$\frac{w_{0}}{\tilde{w}_{0}}=\frac{z_{\text{R}}^{2}}{\tilde{z}_{\text{R}}^{2}}=b^{2}.$ (71)
Then we obtain
$w(z+z_{0})$ $=w_{0}\sqrt{1+\left(\frac{z}{z_{\text{R}}}\right)^{2}}$ (72)
$=b^{2}\tilde{w}_{0}\sqrt{1+\left(\frac{z}{b\tilde{z}_{\text{R}}}\right)^{2}}$ (73)
$=b^{2}\tilde{w}\left(\frac{z}{b}\right).$ (74)
@@
Thus, the new beam $w$ can be obtained from the reference beam $\tilde{w}$ by a scaling factor $b^{2}$ along the transverse coordinates ($x$ and $y$) and by a scaling factor $b$ along the longitudinal propagation coordinate $z$ [noting that the beam is increased in size for $b>1$ along transverse as well as longitudinal coordinates despite the factor $b$ appearing once in the numerator and once in the denominator of Eq. (74)]. The result is then shifted by $-z_{0}$, i.e., by $+\operatorname{Re}q$, along the positive propagation direction.
@@
For representing finite-length beams, one can “clip” the resulting object appropriately at its end positions such that only the relevant section (e.g., between two mirrors) is displayed.
5.2 Three-dimensional beam model
@@
Now that we have seen in Section 5.1 how any Gaussian laser beam can be obtained from one given reference shape, we discuss how such a reference shape can be prepared efficiently. In three-dimensional computer graphics, solid objects are generally modeled as polygon meshes, i.e., a set of points in three-dimensional space (called “vertices”) that are connected by straight lines (“edges”) that in turn form small planes (“faces”). The collection of vertices and lines is sought such that the resulting set of faces approximates the real surface of the given object.
@@
In practice, striving for good performance in real-time rendering applications, one seeks to keep the number of vertices as small as possible to reduce calculation time, yet still capturing the essential shape of the underlying object as accurately as possible. In the present case, we desire to find the minimum number of vertices and their positions that can be used to represent a Gaussian laser beam. The challenge will be that the beam radius of a Gaussian evolves as a hyperbola along the propagation direction as given by Eq. (61) and as a circle around the circumference with radius $w$, while the edges of the mesh are straight line segments. For optimizing the number and distribution of vertices, we introduce a local relative length error, $\varepsilon$, defined as the relative length difference between a curve segment following the real Gaussian shape and a straight edge, connecting the same two neighboring vertices. Then we demand this relative error to be below a certain threshold for all the edges in the mesh, and we seek a vertex distribution that minimizes the number of required vertices for a user-specified error.
We derive in Section LABEL:SISecOptimalVertexDistribution of the SI the minimum number of vertices along the propagation direction for any Gaussian beam, independent of its beam parameters, and find that just five vertices along the propagation direction are sufficient to ascertain a relative maximum error of $\varepsilon=1\,\mathrm{\%}$. This signifies that in many cases, Gaussian beams are well approximated by straight lines at long distances, i.e., the asymptotes of the hyperbolic divergence that is found for $z\gg\tilde{z}_{\text{R}}$ in Eq. (61), and that for the curving sections (between the Rayleigh points) few additional vertices are sufficient. Around the circumference, we find that $13$ vertices are required for $\varepsilon=1\,\mathrm{\%}$, signifying that, compared to the longitudinal direction, more straight edges are required to approximate a circle. Altogether, thus, one needs a minimum of $5\times 13=65$ vertices. Requiring an error of $\varepsilon=0.5\,\mathrm{\%}$ changes these numbers to $7\times 18=126$.
In practice, it may not be practical to use so few vertices along the longitudinal direction because then the resulting mesh might be much larger than the required beam segment between two optical elements. In that case, there would be no mesh vertices at all in the required spatial region. While it is still possible, using suitable rendering agorithms with clipping, to display objects whose vertices are all located outside of the visible region, it may become preferable to define the object with closer-spaced vertices. In particular, the mesh vertices may be required not only for a faithful representation of an (infinitely long) Gaussian beam, but they are also used to terminate the graphical representation when the laser intersects with an object. For that reason, it is helpful to adopt a smaller-step longitudinal mesh spacing in practice by selecting an appropriate maximum step size between any two vertices along the propagation direction, i.e., by selecting an appropriate $\Delta z_{\text{max}}$ in Eq. (LABEL:SIEqOptimalLongitudinalMeshSpacing). The present analysis assures, however, that the relative local error $\varepsilon$ will always be smaller than the given error threshold, and thus the graphical output will approximate the real beam shape well.
5.3 Laser color
@@
When a laser pulse is scattered off a surface, an observer can perceive the spot with a certain color and a certain luminance (“brightness”) that depends on the pulse spectrum, energy, beam radius, and surface properties. For a realistic visualization of such a cross section or of the full laser-beam mesh discussed in Section 5.1 and Section 5.2, we should assign a corresponding computer-graphics color code. For this purpose, we adapt procedures from computer graphics visualization.
@@
In Section LABEL:SISecColorPerception of the SI, we use a “standard observer” model of average human color perception as defined by the “Commission Internationale de l’Éclairage” (CIE). We use the XYZ standard in its most recent implementation [48] and employ the “2-deg XYZ color matching functions transformed from the CIE (2006) 2-deg LMS cone fundamentals” to get the XYZ color values for the given laser spectrum. This color code is subsequently transformed to an RGB representation that can be displayed in computer graphics. In addition, one can use the “alpha” channel A in the RGBA system (typically representing transparency or opacity) to define luminance, as shown in Eq. (LABEL:SIEqLuminance). This luminance is proportional to the pulse-energy scaling factor $S$, meaning a laser beam with higher energy will appear brighter.
@@ 6 Simulation results
Having introduced our model for linear and nonlinear optics simulations, we now show exemplary simulation results. For that purpose, we consider a setup for a well-known pulse-characterization method, SHG frequency-resolved optical gating (FROG) [49, 26]. The setup as shown in Fig. 8(a) is based on a Mach–Zehnder-type interferometer. Note that, to achieve perfect interferometer alignment, the GOEs shown in Fig. 8(a) were placed programmatically by setting numerical position values rather than by placing them “by hand” in VR. The laser emitter in the top-left corner of Fig. 8(a) generates Gaussian pulses with a center wavelength of 800 nm and $\tau_{\text{p}}=15$ fs duration [intensity full width at half maximum as in Eq. (LABEL:SIEqNonNormalizedTemporalEnvelope)]. After passing an idealized beam splitter without dispersion, the reflected beam with wave vector $\mathbf{k}_{1}$ is delayed using a delay stage which can be controlled “remotely” through a virtual computer interface within the VR environment. The transmitted beam with wave vector $\mathbf{k}_{2}$ is noncollinearly overlapped with the reflected beam in a GOE that has the material parameters of a $\beta$-barium borate (BBO) crystal. Under the assumption of perfect phase matching, the second-order response of the BBO was modeled to be “complete,” i.e., the second harmonics of the two individual beams in $2\mathbf{k}_{1}$ and $2\mathbf{k}_{2}$ direction are generated in addition to the “collaborative” SHG in $\mathbf{k}_{1}+\mathbf{k}_{2}$ direction [Fig. 8(b)]. The latter was detected by a spectrometer. In the scan procedure in VR, the delay stage was incremented from $-100$ fs to $+100$ fs in steps of $0.5$ fs. Rather than directly calculating the resulting FROG trace from an idealized “FROG signal equation,” the data was acquired analogously to reality, i.e., after moving the delay stage to a new step position, the SHG spectrum resulting from the interference of $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ in the BBO was calculated according to Section 4.3 before moving to the next step position.
The entire scan took $22\,\mathrm{s}$, and the consistently high and stable frame rate allowed real-time monitoring of changes in the fringe pattern detected by the spectrometer at each scanning step. The thus-obtained “experimental” FROG trace shown in Fig. 8(c) was evaluated using a commercial FROG analysis program [50], yielding the reconstruction results shown in Fig. 8(d–f).
@@ 7 Conclusion
In the present work, we have described the development of a model for real-time simulation of linear and nonlinear optical phenomena and spectroscopy. The term “real time” means that we endeavor to find such a level of approximation that allows calculation times of complete experimental configurations on the order of $10\,\mathrm{ms}$ on consumer-grade hardware. We chose to implement a mixture of geometrical optics and Gaussian wave optics because the former offers computational speed and simplicity, while the latter provides the correct evolution of beam radius and beam curvature during propagation. A finite beam radius is in turn important to describe effects of beam overlap in interference phenomena and nonlinear signal generation. A summary of features and limitations is given in tabular form in Section LABEL:SISecFeaturesAndLimitations of the SI.
In particular, we provided expressions that can be used to computationally process frequency- and time-sampled electric field evolutions of femtosecond laser pulses. We removed fast phase oscillations due to the carrier frequency (from the time-domain field) and due to propagation (from the frequency-domain field) to facilitate numerical stability. In addition, a suitable normalization condition separates the pulse energy and the amplitude changes that occur due to beam-radius evolution.
@@
Considering multi-beam linear interference, we obtained a contrast visibility factor that takes into account the individual parameters of each incident beam.
Thus, we treated the case that an arbitrary number of laser beams – which arrive from different directions, are laterally displaced, and have different individual amplitudes and beam radii – are superimposed on a detector plane. While not resolving the detailed spatial interference pattern, we obtained integrated results as measured by a spectrometer or power meter.
Laser pulses are modified by passing through GOEs. In particular, we considered the response-function formalism to describe linear non-resonant dispersion, resonant absorption (using a Franck–Condon model), and second-order non-resonant response. The latter allowed inclusion of sum-frequency and second-harmonic generation. While such processes had been treated extensively in the scientific literature, we derived simplified approximative results that nevertheless take into account the Gaussian beam overlap between generating pulses and the Gaussian beam curvatures (i.e., focusing properties), and how both properties translate to the nonlinearly generated signal beam.
We derived general spatial properties of Gaussian laser beams that are relevant for graphical representation. In particular, we found that the spatial envelope of any Gaussian beam can be visualized, at a $1\,\mathrm{\%}$ error level, in three-dimensional space using a polygonal mesh with just five vertices connected by straight edges along the propagation direction to describe the hyperbolic beam evolution and 13 vertices along the circumference to approximate the circular symmetry. Such a rendered object can then be scaled to represent Gaussian beams with any other spatial parameters, and we derived the scaling factors. For color visualization, we also obtained the relation between the physical laser spectrum and the computer-graphics RGB color system.
@@
Data Availability Statement The data that support the findings of this study are openly available at the following DOI: [link will be inserted upon publication].

### v2

@@ 5 Graphical laser beam representation
Simulating linear and nonlinear optical phenomena leads to laser beams propagating through space with parameters as determined in Section 2 and Section 4. If one only seeks numerical results, then the calculation is already complete because the obtained parameters provide a full characterization. Often it is desired, however, to represent the laser beams graphically without compromising the real-time simulation. The details of graphical visualization of laser beams will depend on the particular software implementation, which is beyond the scope of the present work. Instead, we discuss some general points that are relevant for any such three-dimensional modeling that is sufficiently fast and realistic.
@@
In Section LABEL:SISecGaussianBeamScaling of the SI, we show that any Gaussian beam can be derived from the general shape of a “reference beam” by suitable scaling. Thus, it is not necessary to re-calculate, point by point, the spatial envelopes of a beam for any new set of beam parameters, but one can simply generate a Gaussian model once, and then scale it accordingly. Let us assume we have available a representation of a Gaussian “reference beam” that is given by its beam radius $\tilde{w}(z)$ at position $z$ along the propagation direction. From that reference, we can obtain the representation of any new Gaussian beam, $w(z)$, according to Eq. (LABEL:SIEqMeshTransformation),
$w(z+z_{0})=b^{2}\tilde{w}\left(\frac{z}{b}\right),$ (15)
@@
by applying a scaling factor $b^{2}$ along the transverse coordinates ($x$ and $y$) and a scaling factor $b$ along the longitudinal propagation coordinate $z$. The result is then shifted by $-z_{0}$, along the positive propagation direction, to obtain the correct beam-waist position. The scaling factor
$b=\frac{z_{\text{R}}}{\tilde{z}_{\text{R}}}$ (16)
@@
is the ratio of the Rayleigh lengths of new beam and reference beam [Eq. (LABEL:SIEqScalingFactorRayleighLengths)]. For representing finite-length beams, one can “clip” the resulting object appropriately at its end positions such that only the relevant section (e.g., between two mirrors) is displayed.
@@
We now discuss how such a reference shape can be prepared efficiently. In three-dimensional computer graphics, solid objects are generally modeled as polygon meshes, i.e., a set of points (“vertices”) in three-dimensional space that are connected by straight lines (“edges”) that in turn form small planes (“faces”). The collection of vertices and edges is sought such that the resulting set of faces approximates the real surface of the given object.
@@
In practice, striving for good performance in real-time rendering applications, one seeks to keep the number of vertices as small as possible to reduce calculation time, yet still capturing the essential shape of the underlying object as accurately as possible. In Section LABEL:SISecOptimalVertexDistribution of the SI, we find the minimum number of vertices and their positions that can be used to represent a Gaussian laser beam for a given maximum allowed error. That error is defined as the relative length difference between a curve segment following the real Gaussian shape and a straight edge, connecting the same two neighboring vertices. In particular, we found that the spatial envelope of any Gaussian beam can be visualized, at a $1\,\mathrm{\%}$ error level, in three-dimensional space using a polygonal mesh with just five vertices connected by straight edges along the propagation direction to describe the hyperbolic beam evolution and $13$ vertices along the circumference to approximate the circular symmetry.
@@
Lastly, let us discuss the perceived color of a laser beam that is relevant in realistic graphical representations. When a laser pulse is scattered off a surface, an observer can perceive the spot with a certain color and a certain luminance (“brightness”) that depends on the pulse spectrum, energy, beam radius, and surface properties. For a realistic visualization of such a cross section or of the full laser-beam mesh discussed in Sections LABEL:SISecGaussianBeamScaling and LABEL:SISecOptimalVertexDistribution of the SI, we assign a corresponding computer-graphics color code.
@@
In Section LABEL:SISecColorPerception of the SI, we use a “standard observer” model of average human color perception as defined by the “Commission Internationale de l’Éclairage” (CIE). We use the XYZ standard in its most recent implementation [36] to get the XYZ color values for the given laser spectrum. This color code is subsequently transformed to an RGB representation that can be displayed in computer graphics. In addition, one can use the “alpha” channel A in the RGBA system (typically representing transparency or opacity) to define luminance, as shown in Eq. (LABEL:SIEqLuminance). This luminance is proportional to the pulse-energy scaling factor $S$, meaning a laser beam with higher energy will appear brighter.
@@ 6 Simulation results
Having introduced our model for linear and nonlinear optics simulations, we now show exemplary simulation results. For that purpose, we consider a setup for a well-known pulse-characterization method, second-harmonic generation (SHG) frequency-resolved optical gating (FROG) [37, 26]. The setup as shown in Fig. 1(a) is based on a Mach–Zehnder-type interferometer. Note that, to achieve perfect interferometer alignment, the GOEs shown in Fig. 1(a) were placed programmatically by setting numerical position values rather than by placing them “by hand” in VR. The laser emitter in the top-left corner of Fig. 1(a) generates Gaussian pulses with a center wavelength of 800 nm and $\tau_{\text{p}}=15$ fs duration [intensity full width at half maximum as in Eq. (LABEL:SIEqNonNormalizedTemporalEnvelope)]. After passing an idealized beam splitter without dispersion, the reflected beam with wave vector $\mathbf{k}_{1}$ is delayed using a delay stage which can be controlled “remotely” through a virtual computer interface within the VR environment. The transmitted beam with wave vector $\mathbf{k}_{2}$ is noncollinearly overlapped with the reflected beam in a GOE that has the material parameters of a $\beta$-barium borate (BBO) crystal. Under the assumption of perfect phase matching, the second-order response of the BBO was modeled to be “complete,” i.e., the second harmonics of the two individual beams in $2\mathbf{k}_{1}$ and $2\mathbf{k}_{2}$ direction are generated in addition to the “collaborative” SHG in $\mathbf{k}_{1}+\mathbf{k}_{2}$ direction [Fig. 1(b)]. The latter was detected by a spectrometer. In the scan procedure in VR, the delay stage was incremented from $-100$ fs to $+100$ fs in steps of $0.2$ fs. Rather than directly calculating the resulting FROG trace from an idealized “FROG signal equation,” the data was acquired analogously to reality, i.e., after moving the delay stage to a new step position, the SHG spectrum resulting from the interference of $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ in the BBO was calculated according to Section 4.2 before moving to the next step position.
The entire scan took $45\,\mathrm{s}$, and the consistently high and stable frame rate allowed real-time monitoring of changes in the fringe pattern detected by the spectrometer at each scanning step. The thus-obtained “experimental” FROG trace shown in Fig. 1(c) was evaluated using a commercial FROG analysis program [38], yielding the reconstruction results shown in Fig. 1(d–f).
@@ 7 Conclusion
In the present work, we have described the development of a model for real-time simulation of linear and nonlinear optical phenomena and spectroscopy. The term “real time” means that we endeavor to find such a level of approximation that allows calculation times of complete experimental configurations on the order of $8\,\mathrm{ms}$ on consumer-grade hardware. We chose to implement a mixture of geometrical optics and Gaussian wave optics because the former offers computational speed and simplicity, while the latter provides the correct evolution of beam radius and beam curvature during propagation. A finite beam radius is in turn important to describe effects of beam overlap in interference phenomena and nonlinear signal generation.
We summarize in Table 1 basic features and consequences of the physical model of the present work. This list is not meant to be exhaustive but rather to capture essential points for a quick overview.
Table 1: Features and limitations of physical model.
Property Choice Consequence
Beam profile Gaussian Finite beam overlap included
Beam quality $M^{2}$ factor Real beam divergence approximated
Beam symmetry Radially symmetric No astigmatism
Beam clipping No diffraction Geometrical effects included; no gratings
Beam focusing Ideal No chromatic or spherical aberration
Polarization direction Scalar electric fields No birefringence; no polarization
Electric field Discrete sampling Spatial–temporal interference included
First-order response Response function Dispersion and absorption included
Second-order response Instantaneous Pulse-shape effects included
We provided expressions that can be used to computationally process frequency- and time-sampled electric field evolutions of femtosecond laser pulses. We removed fast phase oscillations due to the carrier frequency (from the time-domain field) and due to propagation (from the frequency-domain field) to facilitate numerical stability. In addition, we separated the pulse energy from the amplitude changes that occur due to beam-radius evolution.
@@
Considering multi-beam linear interference, we obtained a contrast visibility factor that takes into account the individual parameters of each incident beam. Thus, we treated the case that an arbitrary number of laser beams – which arrive from different directions, are laterally displaced, and have different individual amplitudes and beam radii – are superimposed on a detector plane. While not resolving the detailed spatial interference pattern, we obtained integrated results as measured by a spectrometer or power meter.
Laser pulses are modified by passing through GOEs. We used the response-function formalism to describe linear non-resonant dispersion, resonant absorption (using a Franck–Condon model), and second-order non-resonant response. The latter allowed inclusion of sum-frequency and second-harmonic generation. While such processes had been treated extensively in the literature, we derived simplified approximate results that nevertheless take into account the Gaussian beam overlap between generating pulses and the Gaussian beam curvatures (i.e., focusing properties), and how both properties translate to the nonlinearly generated signal beam.
We derived spatial properties of Gaussian laser beams for graphical representation. We found the minimum number of vertices of a polygon mesh that keeps the local relative error below a given threshold. Such an object can be scaled to represent Gaussian beams with any other spatial parameters, and we derived the scaling factors. For color visualization, we also obtained the relation between the physical laser spectrum and the computer-graphics RGB(A) color system.
@@
[…]
Author contributions T.B. derived the equations, coded the physical model, and wrote the paper. S.M. coded didactic missions, carried out and analyzed the simulations, and prepared the figures. A.M. coded the femtoPro core software and framework. T.B. and S.v.M. supervised the project. All authors discussed the results and edited the manuscript.
@@
Data Availability Statement The data that support the findings of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.15535423 [39].